OpenStudy (chris215):

Find all solutions to the equation in the interval [0, 2π). cos x = sin 2x

OpenStudy (anonymous):

sin2x=2sinxcosx so: cosx=2sinxcosx or 1=2sinx or sinx=1/2 this is true for x=?

OpenStudy (jdoe0001):

\(\bf cos(x)=sin(2x)\implies cos(x)=2sin(x)cos(x) \\ \quad \\ \cfrac{\cancel{cos(x)}}{\cancel{cos(x)}}=2sin(x)\implies \cfrac{1}{2}=sin(x) \\ \quad \\ sin^{-1}\left( \cfrac{1}{2} \right)=sin^{-1}[sin(x)]\implies ?\)

OpenStudy (chris215):

i got 0, pi/6, 5pi/6, pi

OpenStudy (jdoe0001):

hmmm

OpenStudy (xapproachesinfinity):

there is a mistake in the above work \[\cos x=2\sin x\cos x \Longrightarrow \cos x-2\sin x \cos x=0\] \[\cos x(1-2\sin x)=0 \Longrightarrow \cos x=0 ~~or~~\sin x= 1/2\]

OpenStudy (jdoe0001):

you're correct actually the 0 is part of the angles

OpenStudy (jdoe0001):

right.... should have taken common factor and that'd give a 0 value for the angles

OpenStudy (xapproachesinfinity):

cancellation with cos kills some solutions because it assumes cosx is not zero

OpenStudy (jdoe0001):

i got 0, pi/6, 5pi/6, pi \(\huge \checkmark\)

OpenStudy (xapproachesinfinity):

should be good:)

OpenStudy (chris215):

thanks guys :)

OpenStudy (jdoe0001):

yw

OpenStudy (xapproachesinfinity):

np